A toy car having mass m = 1.10 kg

 collides inelastically with a toy train of mass M = 3.80 kg.

 Before the collision, the toy train is moving in the positive x-direction with a velocity of Vi = 2.45 m/s

 and the toy car is also moving in the positive x-direction with a velocity of vi = 4.95 m/s.

 Immediately after the collision, the toy car is observed moving in the positive x-direction with a velocity of 2.25 m/s.

(a) Determine Vf, the final velocity of the toy train.

 



(b) Determine the change ΔKE

 in the total kinetic energy. Assume friction and the rotation of the wheels are not important so that they do not affect ΔKE.



1
Expert's answer
2022-06-13T08:23:11-0400

(a) The law of conservation of momentum says

"mv_i+MV_i=mv_f+MV_f"

So

"V_f=\frac{m}{M}(v_i-v_f)+V_i\\
=\frac{1.10}{3.80}(4.95-2.25)+2.45=3.23\:\rm m/s"

(b)

"\Delta KE=(\frac{mv_i^2}{2}+\frac{MV_i^2}{2})-(\frac{mv_f^2}{2}+\frac{MV_f^2}{2})"

"=(\frac{1.10*4.95^2}{2}+\frac{3.80*2.45^2}{2})-(\frac{1.10*2.25^2}{2}+\frac{3.80*3.23^2}{2})"

"=2.27\:\rm J"